OpenStudy (johnnydicamillo):

find g'(x) given g(x) = e^x/1-5x

OpenStudy (johnnydicamillo):

\[g(x) = \frac{ e^x }{ 1-5x }\]

OpenStudy (johnnydicamillo):

quotient rule?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

you should get e^x/(1-5x)^2. If you need additional help I can help you contact me at mathutor101@gmail.com.

OpenStudy (johnnydicamillo):

so \[e^x * 1-5x - -5 * e^x\]

OpenStudy (johnnydicamillo):

is that right so far

OpenStudy (zarkon):

add parentheses

OpenStudy (johnnydicamillo):

\[(e^x*1-5x)-(-5*e^x)\]

OpenStudy (zarkon):

no

OpenStudy (zarkon):

\[e^x(1-5x)-(-5e^x)\]

OpenStudy (johnnydicamillo):

Okay can I simplify ?

OpenStudy (zarkon):

sure...factor out \(e^x\)

OpenStudy (johnnydicamillo):

(1-5x)-(-5) = \[(1-5x)+5\]

OpenStudy (zarkon):

\[e^x(1-5x)-(-5e^x)=e^x(1-5x+5)=e^x(6-5x)\]

OpenStudy (zarkon):

that is just the numerator

OpenStudy (johnnydicamillo):

okay, now how do I find the denominator, I am so lost.

OpenStudy (zarkon):

\[\left(\frac{N(x)}{D(x)}\right)'=\frac{D(x)N'(x)-N(x)D'(x)}{D(x)^2}\]

OpenStudy (zarkon):

divide by the denominator squared

OpenStudy (johnnydicamillo):

\[\frac{ e^x(6-5x) }{ (1-5x)^2 }\]

OpenStudy (zarkon):

yes

OpenStudy (johnnydicamillo):

Is that all there is to it? Or is there more?

OpenStudy (zarkon):

that is it

OpenStudy (johnnydicamillo):

Thank You!

OpenStudy (zarkon):

No problem